# Extragradient methods with complexity guarantees for hierarchical variational inequalities

### v1

@@ 1 Introduction
Solving hierarchical equilibrium problems is an increasingly active area in mathematical programming that consists in finding a solution to an outer variational inequality over the solutions of an inner variational inequality. In this paper we consider the following problem. Let $(\mathcal{Z},\langle\cdot,\cdot\rangle)$ be a real Hilbert space, potentially infinite-dimensional. We consider the problem of solving the hierarchical equilibrium problem, in which a hemi-variational inequality ($\operatorname{HVI}$) is solved over the solution set of another mixed variational inequality. Specifically, the class of equilibrium problems we consider is of the form
 $\text{ Find }z^{\ast}\in\mathcal{S}_{2}\text{ s.t. }\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle+g_{1}(z)-g_{1}(z^{\ast})\geq 0\qquad\forall z\in\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$ (P)
where ${\mathsf{F}}_{1},{\mathsf{F}}_{2}:\mathcal{Z}\to\mathcal{Z}$ are monotone operators and $g_{1},g_{2}$ are proper, convex, and lower-semicontinuous functions. We refer to the set $\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$ (assumed to be non-empty) as the solution set of the lower-level problem. In the same vein, we will call $\mathcal{S}_{1}$ the set of solutions of (P). Obviously, these sets are nested in the sense $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$. Problem (P) is the formulation of a system of nested $\operatorname{HVI}$’s (also known as variational inequalities of the second kind). These families of variational problems are prominent in mechanics [goeleven2013variational] and optimal control [Barbu93, de2011optimal]. In finite dimensions, it has become a useful model template to study a plethora of problems in mathematical programming, machine learning, operations research, game theory, and signal processing [malitsky2020golden, InertialFBF, Facchinei:2014aa]. Some motivating examples are given below.
@@ 1.1.1 Simple bilevel optimization
This model problem has been studied in several very recent papers. Important references include [merchav2023convex], who approach this problem via a disciplined proximal splitting ansatz, as well as [AminiYous19, DoronShtern23, yousefian2021bilevel, merchav2024fast, boct2025accelerating].
@@ 1.1.2 Hierarchical variational inequalities
If $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{K}}$ for a non-empty, closed convex set $\mathcal{K}\subseteq\mathcal{Z}$, problem (P) reduces to
@@
This is the problem investigated in the recent papers [samadi_improved_2025, marschner2025tikhonov] and [alves2025inertial].
1.1.3 Equilibrium selection in Nash games
Selecting solutions from a given variational inequality problem is another important application of the model template (P). If ${\mathsf{F}}_{1}(x)=\nabla f_{1}(x)$, then we recover the optimal equilibrium selection problem as
$\min_{x\in\mathcal{S}}\{f_{1}(x)+g_{1}(x)\},\qquad\mathcal{S}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$
[…]
@@ 1.1.4 Hierarchical jointly convex generalized Nash equilibrium problem
In very recent work, [LamSagSIOPT25] studied a projection-based method for solving a challenging class of jointly convex generalized Nash equilibrium problems (GNEPs) featuring hierarchy and non-smooth data. We first show that their model is a special case of the problem (P).
The lower level Nash equilibrium problem.
@@
The data of this game problem enjoy the following assumptions:
@@
    The game’s pseudogradient ${\mathsf{F}}_{2}(y)=(\nabla_{y^{1}}h^{\ell}_{1}(y),\ldots,\nabla_{y^{N}}h^{\ell}_{N}(y))$ is monotone on $\mathcal{Y}\triangleq\prod_{\nu=1}^{N}\mathcal{Y}^{\nu}$;
@@
    $\varphi_{\nu}^{\ell}(\cdot)$ is proper closed convex with compact domain $\operatorname{dom}(\varphi^{\ell}_{\nu})\triangleq\mathcal{A}_{\nu}\subset\mathcal{Y}_{\nu}$.
We can characterize equilibria of the Nash game $\Gamma_{\rm low}=\{h^{\ell}_{\nu},\varphi_{\nu}^{\ell}\}_{1\leq\nu\leq N}$ as solutions to the mixed variational inequality $\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$, where $g_{2}(y^{1},\ldots,y^{N})\triangleq\sum_{\nu=1}^{N}\varphi^{\ell}_{\nu}(y^{\nu}).$ The assumptions on the game problem ensure that $\operatorname{dom}(g_{2})$ is a compact subset of the separable Hilbert space $\mathcal{Y}$. Furthermore, $\mathcal{S}_{2}$ is nonempty, convex, and compact.
The upper level GNEP.
The upper level problem is the Nash game with joint coupling constraint represented by the equilibrium set $\mathcal{S}_{2}$. Let $\mu\in\{1,\ldots,M\}$ denote the labels of upper-level players. Each upper-level player aims to solve the constrained optimization problem
@@
Similar to the upper level NEP, we impose the following assumptions on the problem data:
  1. (d)
@@
  2. (e)
    The operator ${\mathsf{F}}_{1}(x)=(\nabla_{x^{1}}h^{u}_{1}(x),\ldots,\nabla_{x^{N}}h^{u}_{N}(x))$ is monotone on $\mathcal{X}\triangleq\prod_{\nu=1}^{N}\mathcal{X}^{\nu}$;
@@
The decision variables $x^{\mu}$ correspond to blocks $(y^{\nu})_{\nu\in\mathcal{N}_{\mu}}$, in which $\mathcal{N}_{\mu}\subseteq\{1,\ldots,N\}$ represent the lower-level players controlled by upper-level player $\mu$. The blocks can well be overlapping, meaning that individual lower level players can have multiple upper level players. In particular, $\mathcal{X}^{\mu}=\prod_{\nu\in\mathcal{N}_{\mu}}\mathcal{Y}^{\nu}$. The entire hierarchical equilibrium problem thus belongs to our template (P), with $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})$, where $g_{1}(x^{1},\ldots,x^{M})\triangleq\sum_{\mu=1}^{M}\varphi^{u}_{\mu}(x^{\mu}).$
@@ 1.2 Related works
This paper belongs to a subclass of bilevel problems in which a single decision variable is to be designed in order to optimize a function over the set of minimizerrs of another function. To emphasize this special character of such problems are also known as simple bilevel problems [Dempe:2021aa]. Starting with the seminal contribution of [Solodov:2007aa] and [Cabot2005], Tikhonov regularization has become one of the dominant paradigms in the numerical solution of simple bilevel optimization problems. In the context of the present paper, this means we have to consider the iteration-dependent operators
${\mathsf{V}}_{k}\triangleq{\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1},\qquad G_{k}\triangleq g_{2}+\sigma_{k}g_{1},$
@@
where $k$ is the iteration counter and $(\sigma_{k})_{k}$ is a well-chosen sequence. Over the past years, a quite significant list of publications have appeared proving rates of convergence with respect to the inner (lower-level) and the outer (upper-level) problem. Specifically, building on the generalized Tikhonov regularization framework of [Solodov:2007aa], the first results in this direction have been produced in [beck2014first, SabSht17]. [merchav2023convex] develop a proximal subgradient method for convex composite models in the upper and lower level problem. Adaptive versions of a related proximal-gradient strategy are studied in [latafat2025convergence]. Accelerated simple bilevel optimization in the convex non-smooth composite setting have recently been derived in [merchav2024fast]. Beyond proximal-gradient based methods, projection-free schemes are developed in [DoronShtern23] and [giang2024projection]. Moving beyond optimization, there exists a significant literature, rooted in the numerical resolution of inverse problems and signal processing, which studied solution methods for selecting a solution to a variational inequality given a pre-defined selection criterion [ogura2002non, xu2003convergence, marino2011explicit, ono2014hierarchical]. This is known as the equilibrium selection problem and is in fact a special case of the general family of Mathematical Programs under equilibrium constraints (MPEC) [luo1996mathematical, Hintermuller:2014aa]. All these papers focus on asymptotic convergence of the sequence generated by tailor-made numerical algorithms. In this more enhanced setting, the only results on iteration complexity we are aware of are published in the papers [kaushik2023incremental, Kaushik:2021aa, jalilzadeh2024stochastic]. Moving beyond the function-case in the upper-level (equilibrium selection), there exist a few papers on numerical methods for solving nested and hierarchical variational inequalities [van2021regularization, ThongNumerical20, lampariello2022solution]. Among those, only [lampariello2022solution] contains complexity statements.
@@
Recently, the papers [samadi_improved_2025] and [alves2025inertial] have addressed the finite-dimensional variational inequality case, obtained from our model template (P) by imposing $g_{1}=0$ and $g_{2}$ an indicator function over a compact convex set. Both papers derive complexity guarantees for the Korpelevich version of the extragradient method, applied to the Tikhonov regularized variational inequality with the operator ${\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1}$ at iteration $k$. The obtained rates are $O(1/k^{\delta})$ and $O(1/k^{1-\delta})$, where $\delta\in(0,1)$ is a parameter defining the strength of the Tikhonov regularization.
In a very recent work, [LamSagSIOPT25] studied a difficult class of game problems featuring hierarchy and non-smooth local cost functions. We show in Section 6 that their model falls into our general problem framework. Their algorithm is not a full splitting method, and thus not directly comparable to our operator splitting approach.
@@ 1.3 Contributions
  • More general problem formulation: Unlike [samadi_improved_2025, alves2025inertial], our hierarchical equilibrium model (P) contains general convex composite terms $g_{1},g_{2}$. Specifically, these papers treat the special case $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{C}}$ for a compact convex and nonempty set $\mathcal{C}$ in a finite-dimensional euclidean vector space. Our analysis performed in potentially infinite-dimensional real Hilbert spaces, and considers general hemi-variational inequalities ($\operatorname{HVI}$’s). This is class contains a rather large class of variational models, which found significant application in optimal control and mechanics [Facchinei:2014aa, goeleven2013variational]. Besides the gain in modelling, adding these convex and potentially non-smooth functions is an important extension, since it allows us to cover the non-smooth composite convex model when specialized to the potential case, as studied in [merchav2024fast].
@@
  • Enhanced Algorithmic design: The recent papers [samadi_improved_2025, alves2025inertial] approach the numerical resolution of problem (P) via double-call version of the extragradient method. Our approach instead uses the ideas of the optimistic extragradient method [hsieh2019convergence] and thus requires only one operator evaluation per iteration. Within this algorithmic setting, we derive sublinear convergence rates for the lower and upper level problem and improved rates under the strong monotonicity of ${\mathsf{F}}_{1}$. Specifically, using a polynomial regularization sequence $\sigma_{k}=O(k^{-\delta})$ for $\delta\in(0,1)$, we demonstrate upper complexity bounds on the order of $O(k^{-\delta})$ for the gap function associated with the lower level equilibrium problem, and a $O(k^{-(1-\delta)})$ complexity bound for the gap function of the entire hierarchical problem. This matches the existing bounds reported in [samadi_improved_2025, alves2025inertial], but for a much larger class of equilibrium problems, and fewer function evaluations (specifically, instead of 2 just 1 per iteration).
@@
  • Unbounded domains: An important assumption in the complexity analysis in [samadi_improved_2025, alves2025inertial] is compactness of the domains over which the variational inequalities are solved. We abandon this assumption. Instead, our proof builds on ideas originating in the convergence analysis of non-autonomous evolution equations due to [attouch2018asymptotic]. In the potential case, these ideas have already been successfully applied to the study of hierarchical optimization problems [Bot:2014aa, Cabot2005]. In particular, [boct2025accelerating] make the role of this technique very transparent by emphasizing the role of the geometry of the lower level problem in establishing a unified complexity statement for both levels simultaneously. A methodological contribution of this work is to show how these ideas naturally translate to the variational setting. We belief that our proof technique is going to be useful also for other hierarchical equilibrium problems as well.
@@ 2.1 Preliminaries
$\operatorname{\textsf{prox}}_{G}(u)=\operatorname*{argmin}_{z\in\mathcal{Z}}\{G(z)+\frac{1}{2}\norm{z-u}^{2}\}.$
If $x,u\in\mathcal{Z}$, we have
@@
The metric projection onto a nonempty closed convex set $\mathcal{C}\subset\mathcal{Z}$ is defined as $\Pi_{\mathcal{C}}(x)=\operatorname*{argmin}_{y\in\mathcal{C}}\norm{x-y}=\operatorname{\textsf{prox}}_{\operatorname{\iota}_{\mathcal{C}}}(x)$, where the convex indicator function $\operatorname{\iota}_{\mathcal{C}}:\mathcal{Z}\to(-\infty,\infty]$ is defined by
$\operatorname{\iota}_{\mathcal{C}}(x)\triangleq\left\{\begin{array}[]{cc}0&\text{if }x\in\mathcal{C},\\ +\infty&\text{ if }x\notin\mathcal{C}.\end{array}\right$.
We denote also $\operatorname{dist}(x,\mathcal{C})\triangleq\norm{x-\Pi_{\mathcal{C}}(x)}$. The polar cone attached to a closed convex set $\mathcal{C}\subset\mathcal{Z}$ is $\mathcal{C}^{\circ}\triangleq\{z\in\mathcal{Z}|\langle z,x\rangle\leq 0\qquad\forall x\in\mathcal{C}\}$. The conjugate function of $f\in\Gamma_{0}(\mathcal{Z})$ is $f^{\ast}(y)\triangleq\sup_{x\in\operatorname{dom}(f)}\{\langle y,x\rangle-f(x)\}.$ A mapping (operator) ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly-monotone ($\mu\geq 0$) if
@@
A 0-monotone operator is simply called monotone. ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $L$-Lipschitz continuous if
@@
The normal cone of a closed convex set $\mathcal{C}\subset\mathcal{Z}$ at point $z\in\mathcal{C}$ is defined $\operatorname{\mathsf{NC}}_{\mathcal{C}}(z)\triangleq\{\xi\in\mathcal{Z}|\langle\xi,z^{\prime}-z\rangle\leq 0\qquad\forall z^{\prime}\in\mathcal{C}\}$. Under convexity, the tangent cone $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)$ of $\mathcal{C}$ at $x\in\mathcal{C}$ is polar to the normal cone, and we have the explicit expression $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)=\operatorname{cl}\left(\bigcup_{\lambda>0}\frac{\mathcal{C}-x}{\lambda}\right)$. The support function of set $\mathcal{C}$ is defined as $\operatorname{\mathtt{s}}(p|\mathcal{C})\triangleq\sup_{z\in\mathcal{C}}\langle p,z\rangle.$ It follows from these definitions that
Lemma 2.1 (Lemma 5.31, [BauCom16]).
[…]
$a_{k+1}\leq a_{k}-b_{k}+c_{k}.$
Then, $\lim_{k\to\infty}a_{k}$ exists and $\sum_{k\geq 1}b_{k}<\infty$.
2.2 Hemi-variational inequalities
Given a monotone and Lipschitz continuous mapping ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ and a function $g\in\Gamma_{0}(\mathcal{Z})$, the hemi-variational inequality problem $\operatorname{HVI}({\mathsf{F}},g)$ is to find $z^{\ast}\in\mathcal{Z}$ such that

### v2

@@ 1 Introduction
Let $(\mathcal{Z},\langle\cdot,\cdot\rangle)$ be a real separable Hilbert space. We consider the problem of solving the hierarchical equilibrium problem, in which a hemi-variational inequality ($\operatorname{HVI}$), also known as variational inequality of the second kind, is solved over the solution set of another $\operatorname{HVI}$:
 $\text{ Find }z^{\ast}\in\mathcal{S}_{2}\text{ s.t. }\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle+g_{1}(z)-g_{1}(z^{\ast})\geq 0\qquad\forall z\in\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}),$ (P)
where ${\mathsf{F}}_{1},{\mathsf{F}}_{2}:\mathcal{Z}\to\mathcal{Z}$ are monotone operators and $g_{1},g_{2}$ are proper, convex, and lower-semicontinuous functions. We refer to the set $\mathcal{S}_{2}$ (assumed to be non-empty) as the solution set of the lower-level problem. In the same vein, we will call $\mathcal{S}_{1}$ the set of solutions of (P). Obviously, these sets are nested in the sense that $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$. Hierarchical $\operatorname{HVI}$’s as in (P) are prominent in mechanics [28] and optimal control [8, 19]. In finite dimensions, it has become a useful model template to study a plethora of problems in mathematical programming, machine learning, control, game theory, and signal processing [42, 16, 24]. Some motivating examples are given below.
@@ 1.1.1 Simple bilevel optimization
This model has been studied in several very recent papers, e.g., in [48], which approach this problem via a disciplined proximal splitting ansatz, as well as in [3, 21, 65, 47, 13].
@@ 1.1.2 Hierarchical variational inequalities
If $g_{1}=0$ and $g_{2}$ is the convex indicator of a non-empty, closed convex set $\mathcal{K}\subseteq\mathcal{Z}$, problem (P) reduces to
@@
This is a challenging formulation of nested variational inequalities, which has been recently investigated in [57, 46, 2].
@@ 1.1.3 Hierarchical jointly convex generalized Nash equilibrium problem
A very recent work [38] studied a projection method for solving a challenging class of jointly convex generalized Nash equilibrium problems (GNEPs) featuring hierarchy and non-smooth data (see [17] for a stochastic formulation of a similar problem).
The lower-level Nash Equilibrium Problem (NEP)
@@
The data of this game problem enjoys the following assumptions:
@@
    The game’s pseudogradient ${\mathsf{F}}_{2}(y)\triangleq(\nabla_{y^{1}}h^{\ell}_{1}(y),\ldots,\nabla_{y^{N}}h^{\ell}_{N}(y))$ is monotone on the separable Hilbert space $\mathcal{Y}\triangleq\prod_{\nu=1}^{N}\mathcal{Y}^{\nu}$;
@@
    The functions $\varphi_{\nu}^{\ell}(\cdot)$ are proper, closed, and convex, with compact domain.
We can characterize equilibria of the Nash game $\Gamma_{\rm low}\triangleq\{h^{\ell}_{\nu},\varphi_{\nu}^{\ell}\}_{1\leq\nu\leq N}$ as solutions to the $\operatorname{HVI}$ with data ${\mathsf{F}}_{2}$, $g_{2}(y^{1},\ldots,y^{N})\triangleq\sum_{\nu=1}^{N}\varphi^{\ell}_{\nu}(y^{\nu})$ (see e.g. [25]). Hence, $\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$. The above assumptions ensure that $\operatorname{dom}(g_{2})$ is a compact subset of $\mathcal{Y}$. Furthermore, $\mathcal{S}_{2}$ is nonempty, convex, and compact.
The upper level GNEP
The upper-level problem is the Nash game with joint coupling constraint represented by the equilibrium set $\mathcal{S}_{2}$. Let $\mu\in\{1,\ldots,M\}$ denote the labels of upper-level players. Each upper-level player aims to solve the constrained optimization problem
@@
As for the lower-level NEP, we make the following assumptions on the problem data:
  1. (e)
@@
    The operator ${\mathsf{F}}_{1}(x)\triangleq(\nabla_{x^{1}}h^{u}_{1}(x),\ldots,\nabla_{x^{N}}h^{u}_{N}(x))$ is monotone on the product space $\mathcal{X}\triangleq\prod_{\nu=1}^{N}\mathcal{X}^{\nu}$;
  3. (g)
@@
The decision variables $x^{\mu}$ correspond to blocks $(y^{\nu})_{\nu\in\mathcal{N}_{\mu}}$, in which $\mathcal{N}_{\mu}\subseteq\{1,\ldots,N\}$ represent the lower-level players controlled by upper-level player $\mu$. In particular, $\mathcal{X}^{\mu}=\prod_{\nu\in\mathcal{N}_{\mu}}\mathcal{Y}^{\nu}$. The entire hierarchical equilibrium problem thus belongs to our template (P), with $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})$, where $g_{1}(x^{1},\ldots,x^{M})\triangleq\sum_{\mu=1}^{M}\varphi^{u}_{\mu}(x^{\mu}).$ Note that if $M=1$ (i.e., there is only one player on the upper level), then we recover the problem of equilibrium selection in Nash games, which is an important problem in game theory [65, 11].
@@ 1.2 Related works
This paper focuses on a subclass of bilevel problems, usually termed as simple bilevel problems [20], in which a single decision variable is to be designed in order to solve the upper-level problem over the solution set of a lower-level problem. Starting with the seminal contributions [59, 1], Tikhonov regularization has become one of the dominant paradigms in the numerical solution of simple bilevel optimization problems. In our $\operatorname{HVI}$ context, this idea leads to the definition of the data that combines both levels in (P) as
${\mathsf{V}}_{k}\triangleq{\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1},\qquad G_{k}\triangleq g_{2}+\sigma_{k}g_{1},$ (1.1)
@@
where $k$ is the iteration counter and $(\sigma_{k})_{k\geq 1}$ is a well-chosen regularization sequence. Over the past years, quite a significant list of publications appeared, proving rates of convergence with respect to the inner (lower-level) and the outer (upper-level) problem in the setting of hierarchical optimization as in Section 1.1.1. The first results in this direction are [10, 56] for strongly convex upper level functions. [48] develop an alternating proximal subgradient method for convex composite models in the upper and lower level problem. Building on the iterative regularization structure, adaptive versions of a related proximal-gradient strategy are studied in [39] without rates on the upper level problem. [58] and [13] present rates for both upper and lower level problems for the iterative proximal gradient method and its accelerated version, which was first analyzed in [47]. Beyond proximal-gradient based methods, projection-free schemes are developed in [21, 27]. Moving beyond optimization, there exists a significant body of literature, rooted in the numerical resolution of inverse problems and signal processing, that studied solution methods for selecting a solution to a variational inequality given a pre-defined selection criterion [51, 64, 45, 52]. This is known as the equilibrium selection problem, a special case of the general family of mathematical programs under equilibrium constraints [41, 29]. All these papers focus on asymptotic convergence of the sequence generated by tailor-made numerical algorithms. The only results on iteration complexity we are aware of are published in the papers [34, 35, 33]. Moving beyond the function-case in the upper-level (equilibrium selection), there exist a few papers on numerical methods for solving hierarchical VIs [63, 60, 37]. Among those, only [37] contains complexity statements.
@@
Recently, the papers [57] and [2] have addressed the finite-dimensional variational inequality case, obtained from our model template (P) by imposing $g_{1}=0$ and $g_{2}$ to be an indicator function over a compact convex set. Both papers derive complexity guarantees for the Korpelevich extragradient method, applied to the Tikhonov regularized variational inequality with the operator ${\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1}$ at iteration $k$. The obtained rates are $O(1/k^{\delta})$ (lower level) and $O(1/k^{1-\delta})$ (upper level), where $\delta\in(0,1)$ is a parameter defining the strength of the Tikhonov regularization.
@@ 1.3 Contributions
More general problem formulation: Unlike [57, 2], our hierarchical equilibrium model (P) contains general convex composite terms $g_{1},g_{2}$. Specifically, these papers treat the special case $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{C}}$, the indicator function of a compact convex and nonempty set $\mathcal{C}$ in a finite-dimensional Euclidean vector space. Our analysis is performed in potentially infinite-dimensional real Hilbert spaces, and considers general $\operatorname{HVI}$’s. This class contains a rather large set of variational models, which have found significant application in optimal control and mechanics [24, 8, 28]. Besides the gain in modeling, adding these convex and potentially non-smooth functions is an important extension, as it allows treating the non-smooth composite convex model when specialized to the potential case, as studied in [47].
@@
Enhanced Algorithmic design and combined rate statements: The recent papers [57, 2] approach the numerical resolution of problem (P) via double-call version of the extragradient method. Our approach instead builds on the optimistic extragradient method [30], requiring only one operator evaluation per iteration. Within this algorithmic setting, we derive sublinear convergence rates for the lower- and upper-level problem and improved rates under the strong monotonicity of ${\mathsf{F}}_{1}$. Specifically, using a polynomial regularization sequence $\sigma_{k}=O(k^{-\delta})$ for $\delta\in(0,1)$, we demonstrate upper complexity bounds on the order of $O(k^{-\delta})$ for the gap function associated with the lower-level equilibrium problem (feasibility gap), and a $O(k^{-(1-\delta)})$ complexity bound for the gap function of the entire hierarchical problem (optimality gap). These bounds match the existing bounds reported in [57, 2], while applying to a much broader class of equilibrium problems and requiring fewer function evaluations.
@@
Unbounded domains: An important assumption in the complexity analysis in [57, 2] is compactness of the domains over which the variational inequalities are solved. We abandon this assumption. Instead, our proof builds on ideas originating from the asymptotic analysis of non-autonomous evolution equations due to [4]. In the potential case, these ideas have already been successfully applied to the study of hierarchical optimization problems [14, 1]. In particular, [13] make the role of this technique very transparent by emphasizing the importance of the geometry of the lower level problem in establishing a unified complexity statement at both levels simultaneously. A methodological contribution of this work is to show how these ideas naturally translate to the variational setting. We believe that our proof technique is going to be useful for other hierarchical equilibrium problems as well.
The outline of this paper is as follows. Section 2 collects relevant facts on VIs, defining the geometric properties we assume for the solution set of the lower-level problem, and introducing the merit functions we employ in our complexity statements. Section 3 contains the main algorithmic paradigm we employ to prove our rates. Section 4 details the proof of our two main theorems. Section 5 reports important extensions. We establish improved rates under strong monotonicity of the upper-level $\operatorname{HVI}$ and show that the main statements extend to optimistic versions of the forward-backward-forward splitting, enjoying only one proximal operator evaluation per iteration. Section 6 presents numerical experiments demonstrating our results.
@@ 2.1 Preliminaries
For a closed convex set $\mathcal{C}\subset\mathcal{Z}$, the metric projection is $\Pi_{\mathcal{C}}(x)\triangleq\operatorname{\textsf{prox}}_{\operatorname{\iota}_{\mathcal{C}}}(x)$, where the convex indicator function $\operatorname{\iota}_{\mathcal{C}}:\mathcal{Z}\to(-\infty,\infty]$ is defined by $\operatorname{\iota}_{\mathcal{C}}(x)=0$ if $x\in\mathcal{C}$, and $\operatorname{\iota}_{\mathcal{C}}(x)=+\infty$ else. We also denote $\operatorname{dist}(x,\mathcal{C})\triangleq\norm{x-\Pi_{\mathcal{C}}(x)}$. The polar cone attached to a set $\mathcal{C}\subset\mathcal{Z}$ is $\mathcal{C}^{\circ}\triangleq\{z\in\mathcal{Z}|\langle z,x\rangle\leq 0\qquad\forall x\in\mathcal{C}\}$. The normal cone at point $z\in\mathcal{C}$ is defined as $\operatorname{\mathsf{NC}}_{\mathcal{C}}(z)\triangleq\{p\in\mathcal{Z}|\langle p,z^{\prime}-z\rangle\leq 0\qquad\forall z^{\prime}\in\mathcal{C}\}$. For any convex $\mathcal{C}\subset\mathcal{Z}$ (not neccesarily closed), the tangent cone $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)$ of $\mathcal{C}$ at $x\in\mathcal{C}$ is polar to the normal cone, and we have the explicit expression $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)=\operatorname{cl}\left(\bigcup_{\lambda>0}\frac{\mathcal{C}-x}{\lambda}\right)$.
The Fenchel conjugate of $f\in\Gamma_{0}(\mathcal{Z})$ is $f^{\ast}(y)\triangleq\sup_{x\in\operatorname{dom}(f)}\{\langle y,x\rangle-f(x)\}.$ Specifically, given a closed convex set $\mathcal{C}\subset\mathcal{Z}$, the Fenchel conjugate of $\operatorname{\iota}_{\mathcal{C}}$ is the support function of set $\mathcal{C}$ defined as $\operatorname{\mathtt{s}}(p|\mathcal{C})\triangleq\sup_{z\in\mathcal{C}}\langle p,z\rangle.$ Note that
A mapping (operator) ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly-monotone ($\mu\geq 0$) if
@@
A 0-monotone operator is simply called monotone. $\mathsf{F}$} is $L$-Lipschitz continuous if
@@
2.2 Hemi-variational inequalities (HVI’s)
Given a monotone and Lipschitz continuous mapping ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ and a function $g\in\Gamma_{0}(\mathcal{Z})$, the $\operatorname{HVI}$ problem $\operatorname{HVI}({\mathsf{F}},g)$ is to find $z^{\ast}\in\mathcal{Z}$ such that